OpenStudy (anonymous):

help! qudratic formula for: (x+1)^2=2(x-1)^2

OpenStudy (anonymous):

1. square each pair of parenthesis and then foil it out. (x+1)(x+I)= 2(x-1)(x-1) x^2+2x+1=2(x^2-2x+1) 2.distribute the 2 =2x^2-4x+2 3.since quadratics are equal to zero, put everything on one side and combine like terms. 0= -x^2-2x-1+2x^2-4x+2 0= x^2-6x+1 4. That's your quadratic